# The “Little Dark Dot”: Evidence for Self-Interacting Dark Matter in the Strong Lens SDSSJ0946+1006?

### v1

@@ Abstract
Previous studies, based on precise modeling of a gravitationally lensing image, have identified what may be a extremely compact, dark perturber in the well-known lensing system SDSSJ0946+1006 (the “Jackpot”). Its remarkable compactness challenges the standard cold dark matter (CDM) paradigm. In this paper, we explore whether such a compact perturber could be explained as a core-collapsed halo described by the self-interacting dark matter (SIDM) model. Using the isothermal Jeans method, we compute the density profiles of core-collapsed halos across a range of masses. Our comparison with observations indicates that only a core-collapsed halo with a total mass of approximately $10^{11}~{}{\rm M_{\odot}}$ could produce an inner density profile and mass enclosed within 1 kpc that is consistent with observational data. However, such a massive dark matter halo should host a galaxy detectable by prior Hubble imaging, which is not observed. Thus, the core-collapsed SIDM halo model struggles to fully account for the exotic nature of the “little dark dot” in the “Jackpot” lens.
@@ 1 Introduction
The standard theory of cosmic structure formation is built upon the cold dark matter (CDM) model. In this framework, structures grow hierarchically: smaller dark matter halos form first and later merge and accrete material to evolve into larger systems. Galaxies form within these halos and co-evolve with their hosts. A direct prediction of this paradigm is the existence of a large number of low-mass dark matter halos in the universe today.
@@
Detecting these predicted CDM halos is challenging. These halos can host dwarf galaxies, which serve as tracers, but in the distant universe, even if dwarf galaxies are observed, measuring the masses of their host halos is often difficult. Moreover, baryonic feedback processes are expected to suppress star formation in halos smaller than $10^{9}~{}{\rm M_{\odot}}$ (e.g. T. Sawala et al., 2015, 2016; J. S. Bullock & M. Boylan-Kolchin, 2017), leaving many low-mass halos either completely dark or hosting faint galaxies that are hard to detect.
@@
Beyond the Milky Way, strong gravitational lensing provides the only method currently available to directly detect these dark halos and measure their masses and internal structures. Such halos, if present, can perturb lensing systems by introducing anomalies in flux ratios or causing subtle distortions in the lensing images (S. Mao & P. Schneider, 1998; S. Vegetti & L. V. E. Koopmans, 2009). These dark halos may reside within the main lens galaxy or along the line of sight to the background source (R. Li et al., 2016, 2017; G. Despali et al., 2018; D. Gilman et al., 2018; Q. He et al., 2022a, b).
@@
Over the past decade, several candidate dark halos have been identified in high-resolution imaging (S. Vegetti et al., 2010, 2012; Y. D. Hezaveh et al., 2016). Among these, SDSSJ0946+1006 stands out as particularly intriguing. This strong lensing system was first discovered in the SDSS main spectroscopic sample by the SLACS survey and subsequently received deep follow-up observations with the Hubble Space Telescope (HST), accumulating a total exposure time of 2096 seconds (I-band R. Gavazzi et al., 2008). It has become one of the most extensively studied systems, with multiple analyses consistently identifying evidence for a dark perturber (S. Vegetti et al., 2010; Q. Minor et al., 2021; J. W. Nightingale et al., 2024; D. J. Ballard et al., 2024; G. Despali et al., 2024; W. J. R. Enzi et al., 2024). Using pixelised image modeling of the lensing data, S. Vegetti et al. (2010) reported the detection of a dark perturber in this system. Fitting the perturber with a tidally truncated Pseudo-Jaffe profile, characterized by an inner density profile scaling as $\rho\sim r^{-2}$, they estimated a subhalo mass of $3.51\times 10^{9}~{}{\rm M_{\odot}}$.
@@
However, in a CDM universe, pure dark matter halos are expected to follow a Navarro-Frenk-White (NFW) profile (J. F. Navarro et al., 1997), characterized by an inner density profile scaling as $\rho\sim r^{-1}$, which is much shallower than that of a Pseudo-Jaffe profile. G. Despali et al. (2018) demonstrated that fitting an NFW perturber with a Pseudo-Jaffe model could lead to halo mass underestimation by an order of magnitude; they corrected the perturber’s mass to $2.06\times 10^{10}~{}{\rm M_{\odot}}$. Subsequently, Q. Minor et al. (2021) adopted a truncated NFW (tNFW) model, allowing the concentration parameter to vary freely. They found that the perturber is exceptionally compact, with a subhalo mass of $5.01\times 10^{9}~{}{\rm M_{\odot}}$ and an extremely-high concentration of 1560. The inferred parameters become less extreme when “multipoles” are included in the macro model, suggesting deviations from ellipticity in the primary lens. The best-fit values are $M_{200}=3.02\times 10^{10}~{}{\rm M_{\odot}}$ and $c_{200}=70.6$, which still remains at least $5\sigma$ above the median concentration of CDM field halos derived from $N$-body cosmological simulations. These findings have been corroborated by subsequent studies (G. Despali et al., 2024; W. J. R. Enzi et al., 2024).
@@
It is worth noting that the perturber’s compactness cannot be easily explained by assuming it lies along the line of sight rather than in the lens plane. Q. He et al. (2022c) argued that the perturbation induced by a dark halo is strongest when it resides in the lens plane. For a perturber along the line of sight to produce comparable distortions, it would require a higher mass or a steeper density profile. W. J. R. Enzi et al. (2024) allowed the perturber’s redshift to vary in their analysis to the system and concluded that the subhalo is more likely to reside in the lens plane.
@@
The unusual steepness of this halo has prompted consideration of alternative dark matter models, particularly self-interacting dark matter (SIDM), which has been proposed as a potential revision to CDM on small scales (D. N. Spergel & P. J. Steinhardt, 2000; S. Tulin & H.-B. Yu, 2018). In a self-gravitating system, the behavior of dark matter exhibits similarities to that of stars in a globular cluster, and the evolution of an SIDM halo ultimately leads to the “gravothermal catastrophe” (D. Lynden-Bell & R. Wood, 1968): over sufficiently long timescales, the self-interaction-driven dark matter core becomes thermally unstable. This instability leads to gravitational collapse, ultimately re-establishing a cuspy density profile in the halo center (e.g. S. Balberg et al., 2002; S. Balberg & S. L. Shapiro, 2002; J. Pollack et al., 2015; R. Essig et al., 2019; H. Nishikawa et al., 2020). Drawing on the first group-scale $N$-body simulation with large-amplitude, velocity-dependent DM self-interactions, E. O. Nadler et al. (2023) demonstrated that SIDM can produce halos with properties similar to the J0946+1006 perturber. However, comprehensive studies on whether SIDM core-collapsed halos can explain the dark perturber in J0946 remain limited, as most previous analyses have been qualitative rather than quantitative (e.g. E. O. Nadler et al., 2023; W. J. R. Enzi et al., 2024).
@@
In this paper, we systematically investigate the possibility that the “little dark dot” in J0946+1006 is a core-collapsed SIDM halo. We compute the density profiles using an improved version of the isothermal Jeans model that can capture the onset of core-collapse, across a wide range of halo mass, and compare the model predictions with the lensing data. Additionally, we generate mock lensing images with the same configuration as J0946+1006, incorporating a theoretical SIDM halo profile to simulate image distortions. By fitting the distortions, we evaluate the biases introduced during modeling. Our results indicate that an SIDM halo with a mass of $10^{11}~{}{\rm M_{\odot}}$ is required to produce the observed lensing perturbation. However, this raises a new question: Why does not such a massive halo host a prominent visible galaxy?
@@
This paper is organized as follows. Section 2 presents our methodology for modeling SIDM halos (§$\S$ 2.1), simulating their evolutionary processes (§$\S$ 2.2), and computing their lensing properties (§$\S$ 2.3). In Section 3, we compare the results of our SIDM modeling with observational result (§$\S$ 3.1) and evaluate the performance of gravitational lens fitting using the mock lensing image (§$\S$ 3.2). Section 4 presents a discussion of our results, focusing on their implications and limitations, and finally, Section 5 summarizes the principal conclusions of this study and outlines potential directions for future research.
@@
Throughout this paper, we adopt the Planck ${\rm\Lambda CDM}$ cosmology (Planck Collaboration et al., 2016) with $\Omega_{m}=0.307$ and $H_{0}=67.7~{}{\rm km}~{}{\rm s}^{-1}~{}{\rm Mpc}^{-1}$.
@@ 2.1 Solving the density profile of the SIDM halo
Considering its low computational cost and the convenience of being able to model specific galaxies and dark matter halo systems, we use a semi-analytical method called isothermal Jeans method to construct the SIDM halo. This method was first proposed by M. Kaplinghat et al. (2014, 2016) and its efficacy for SIDM density profiles has been intensively tested using hydrodynamical simulations of SIDM (e.g., A. Robertson et al., 2021). Here, we present the basic idea and implementation process of this method. For more detailed information, readers are referred to F. Jiang et al. (2023), where a public code implementation is also provided.
The model is based on the idea that energy transfer through elastic scattering between dark matter particles leads to the formation of an isothermal SIDM core after several interactions over the halo’s evolutionary timescale. The density of this isothermal core $\rho_{\rm iso}$ should be describable by the Jeans equation:
@@
where $\Phi$ is the total gravitational potential, including both dark matter and baryons. Assuming an isotropic ($\beta=0$) and constant velocity dispersion ($\nu=\nu_{0}$) in the core region, we get a simple generic solution of the Jeans equation:
@@
where $\rho_{\rm 0}$ is the central density of the isothermal core and $\Delta\Phi=\Phi(r)-\Phi(0)$. To obtain the expression of $\Phi(r)$, we apply the Poisson equation, which connects the gravitational potential to the matter density:
@@
where $\rho_{\rm b}$ represents the density distribution of the galaxy. The dark matter density (Eq. 2) and the galaxy’s density are then substituted into the Poisson equation (Eq. 3). By applying appropriate variable transformations and setting the boundary conditions, the resulting second-order differential equation is solved to obtain $\rho_{\rm iso}$ as a function of radius. More details for each steps can be found in Appendix A.
For a given set of $\rho_{\rm 0}$ and $\nu_{0}$, we can determine the density profile of the isothermal core of the SIDM halo following the above steps. However, elastic scattering between dark matter particles primarily occurs in the central regions, where the particle density is sufficiently high. In the outer regions of the dark halo, the interaction between particles becomes negligible, and the dark matter distribution still follows the predictions of CDM, i.e. the NFW profile:
@@
After defining $r_{200}$ as the radius within which the the mean enclosed density is 200 times the critical density $\rho_{\rm crit}$, and $M_{200}$ as the mass within $r_{200}$, The scaling parameters $\rho_{s}$ and $r_{s}$ in Eq. 4 can be replaced with the halo mass $M_{200}$ and the concentration parameter $c_{200}=r_{200}/r_{s}$. Now, the SIDM halo is divided into two regions, and we need to identify the boundary that separates them. The stitching radius $r_{1}$ is defined as the radius at which the local scattering rate $\Gamma(r)$, multiplied by the halo’s lifetime $t_{\rm age}$, equals one, marking the transition point where the behavior of the dark halo changes:
@@
Figure 1: The temporal evolution of the central density $\rho_{0}$ (left panel) and velocity dispersion $\nu_{0}$ (right panel) of an SIDM core with $M_{200}=1\times 10^{11}~{}{\rm M_{\odot}}$. Considering the degeneracy between the formation time $t_{\rm age}$ and the scattering cross-section $\sigma_{m}$, we choose their product as the horizontal axis, with time measured in units of the cosmic time of J0946+1006 (11.06 Gyr). The blue, orange, and magenta scatter points represent the evolution for the low-density solution, high-density solution, and mirror high-density solution, respectively. The moments corresponding to $t_{\rm merge}$ and $t_{\rm coll}$ are marked by a green dashed line and a purple dash-dotted line, respectively. The black dots represent the moments we chose to examine the evolution of the SIDM profile, while the red star represents the moment closest to the result of Q. Minor et al. (2021) in the 2D observational space.
Here, $\sigma_{m}$ denotes the effective self-interaction cross section per particle mass (in units of ${\rm cm^{2}/g}$), which we treat as a constant following D. Yang & H.-B. Yu (2022)’s framework. This approximation replaces the velocity-dependent cross section with a single characteristic value for each halo. $\rho_{\rm dm}$ is the dark matter density profile, which can either be set by the inner isothermal core or by the CDM outskirts. $\langle v_{\rm pair}(r)\rangle$ denotes the mean pairwise velocity of particles at radius $r$, and for a Maxwell–Boltzmann velocity distribution with a one-dimensional velocity dispersion profile $\nu(r)$, it is given by $\langle v_{\rm pair}(r)\rangle=\frac{4}{\sqrt{\pi}}\nu(r)$.
@@
Once $r_{1}$ is determined, we can stitch these two regions together by imposing the matching conditions: $\rho_{\rm iso}(r_{1})\approx\rho_{\rm cdm}(r_{1})$ and $M_{\rm iso}(r_{1})\approx M_{\rm cdm}(r_{1})$, and then minimizing the “stitching error” term $\delta^{2}$:
@@
Here, we adopt the outside-in approach (L. Sagunski et al., 2021; A. Robertson et al., 2021). For a CDM halo with known $M_{200}$, $c_{200}$, and the inner galaxy’s density distribution, we first compute the NFW profile for its outskirts. For a given $t_{\rm age}\cdot\sigma_{m}$ value, we determine $r_{1}$ using Eq. 5. With a set of ($\rho_{0}$, $\nu_{0}$) values, we then derive the isothermal core density profile by solving the coupled Jeans and Poisson equations (Eq. 1 and Eq. 3). Finally, we optimize the ($\rho_{0}$, $\nu_{0}$) combination to minimize $\delta^{2}$ (Eq. 6), ensuring a smooth transition between the isothermal core and CDM outskirts at $r_{1}$. In summary, once $M_{200}$, $c_{200}$, $t_{\rm age}$, $\sigma_{m}$, and the galaxy’s density profile are provided, this isothermal Jeans model yields the complete density profile of the SIDM halo.
In Appendix A, we discuss the impact of baryonic matter on SIDM halos within our SIDM model. Given the non-detection of an optical counterpart for J0946+1006 and the upper limit of its luminosity should be no bigger than $5\times 10^{6}~{}{\rm L_{\odot}}$ (S. Vegetti et al., 2010), we restrict our analysis exclusively to the dark matter only situation. Specifically, we exclude the galactic density profile from Eq. 3 and employ the NFW profile alone to derive the density profile of the inner SIDM core.

### v2

@@ Abstract
Previous studies, based on precise modeling of a gravitationally lensing image, have identified what may be an extremely compact, dark perturber in the well-known lensing system SDSS J0946+1006 (the “Jackpot”). Its remarkable compactness challenges the standard cold dark matter (CDM) paradigm. In this paper, we explore whether such a compact perturber could be explained as a core-collapse halo described by the self-interacting dark matter (SIDM) model. Using the isothermal Jeans method, we compute the density profiles of core-collapse halos across a range of masses. Our comparison with observations indicates that a core-collapse halo has an inner density profile and mass enclosed within 1 kpc that fit the data well, but only if the halo has a total mass $\sim 10^{11}~{\rm M_{\odot}}$. While a halo of this mass should host a detectable galaxy, the current observational upper limit on the perturber’s luminosity remains uncertain. Resolving whether or not the data support the presence of a core-collapse SIDM halo therefore requires future deep observations to measure its luminosity.
@@ 1 Introduction
The standard theory of cosmic structure formation is built upon the cold dark matter (CDM) model. In this framework, structures grow hierarchically: smaller dark matter halos form first and later merge and accrete material to evolve into larger systems. Galaxies form within these halos and coevolve with their hosts. A direct prediction of this paradigm is the existence of a large number of low-mass dark matter halos in the Universe today.
@@
Detecting these predicted CDM halos is challenging. These halos can host dwarf galaxies, which serve as tracers, but in the distant Universe, even if dwarf galaxies are observed, measuring the masses of their host halos is often difficult. Moreover, baryonic feedback processes are expected to suppress star formation in halos smaller than $10^{9}~{\rm M_{\odot}}$ (e.g., Sawala2015; Sawala2016; Bullock2017), leaving many low-mass halos either completely dark or hosting faint galaxies that are hard to detect.
@@
Beyond the Milky Way, strong gravitational lensing provides the only method currently available to directly detect these dark halos and measure their masses and internal structures. Such halos, if present, can perturb lensing systems by introducing anomalies in flux ratios or causing subtle distortions in lensing images (Mao1998; Vegetti2009). These dark halos may reside within the main lens galaxy or along the line of sight to the background source (Li_cdmwdm2016; Li2017_los; Despali2018; Gilman2018; He2022_los; He2022_abc).
@@
Over the past decade, several candidate dark halos have been identified in high-resolution imaging (Vegetti_2010; Vegetti2012; Hezaveh2016_alma_subhalo). Among these, SDSS J0946+1006 stands out as particularly intriguing. This strong-lensing system was first discovered in the Sloan Digital Sky Survey main spectroscopic sample by the SLACS survey and subsequently received deep follow-up observations with the Hubble Space Telescope (HST), accumulating a total exposure time of 2096 s ($I$-band; Gavazzi_2008). It has become one of the most extensively studied systems, with multiple analyses consistently identifying evidence for a dark perturber (Vegetti_2010; Minor2021; Ballard_24; Enzi2024; Nightingale_24; Despali2024). Using pixelized image modeling of the lensing data, Vegetti_2010 reported the detection of a dark perturber in this system. Fitting the perturber with a tidally truncated pseudo-Jaffe profile, characterized by an inner density profile scaling as $\rho\sim r^{-2}$, they estimated a subhalo mass of $3.51\times 10^{9}~{\rm M_{\odot}}$.
@@
However, in a CDM Universe, pure dark matter halos are expected to follow a Navarro–Frenk–White (NFW) profile (NFW1997), characterized by an inner density profile scaling as $\rho\sim r^{-1}$, which is much shallower than that of a pseudo-Jaffe profile. Despali2018 demonstrated that fitting an NFW perturber with a pseudo-Jaffe model could lead to halo mass underestimation by an order of magnitude; they corrected the perturber’s mass to $2.06\times 10^{10}~{\rm M_{\odot}}$. Subsequently, Minor2021 adopted a truncated NFW (tNFW) model, allowing the concentration parameter to vary freely. They found that the perturber is exceptionally compact, with a subhalo mass of $5.01\times 10^{9}~{\rm M_{\odot}}$ and an extremely high concentration of 1560. The inferred parameters become less extreme when “multipoles” are included in the macro model, suggesting deviations from ellipticity in the primary lens. The best-fit values are $M_{200}=3.02\times 10^{10}~{\rm M_{\odot}}$ and $c_{200}=70.6$, which still remain at least $5\sigma$ above the median concentration of CDM field halos derived from $N$-body cosmological simulations. These findings have been corroborated by subsequent studies (Enzi2024; Despali2024).
@@
It is worth noting that the perturber’s compactness cannot be easily explained by assuming it lies along the line of sight rather than in the lens plane. He2022_los argued that the perturbation induced by a dark halo is strongest when it resides in the lens plane. For a perturber along the line of sight to produce comparable distortions, it would require a higher mass or a steeper density profile. Enzi2024 allowed the perturber’s redshift to vary in their analysis of the system and concluded that the subhalo is more likely to reside in the lens plane.
@@
The unusual steepness of this halo has prompted consideration of alternative dark matter models, particularly self-interacting dark matter (SIDM), which has been proposed as a potential revision to CDM on small scales (Spergel2000; Tulin&Yu). In a self-gravitating system, the behavior of dark matter exhibits similarities to that of stars in a globular cluster, and the evolution of an SIDM halo ultimately leads to the “gravothermal catastrophe” (Lynden&Bell1968): over sufficiently long timescales, the self-interaction-driven dark matter core becomes thermally unstable. This instability leads to gravitational collapse, ultimately reestablishing a cuspy density profile in the halo center (e.g., Balberg2002PRL; Balberg2002ApJ; Pollack2015; Essig2019; Nishikawa2020). Drawing on the first group-scale $N$-body simulation with large-amplitude, velocity-dependent dark matter self-interactions, Nadler2023 demonstrated that SIDM can produce halos with properties similar to the J0946+1006 perturber. However, comprehensive studies on whether SIDM core-collapse halos can explain the dark perturber in J0946 remain limited, as most previous analyses have been qualitative rather than quantitative (e.g., Nadler2023; Enzi2024).
@@
In this paper, we systematically investigate the possibility that the “little dark dot” in J0946+1006 is a core-collapse SIDM halo. We compute the density profiles using an improved version of the isothermal Jeans model that can capture the onset of core collapse, across a wide range of halo mass and compare the model predictions with the lensing data. Additionally, we generate mock lensing images with the same configuration as J0946+1006, incorporating a theoretical SIDM halo profile to simulate image distortions. By fitting the distortions, we evaluate the biases introduced during modeling. Our results indicate that an SIDM halo with a mass of $10^{11}~{\rm M_{\odot}}$ is required to produce the observed lensing perturbation. However, this raises a new question: why does such a massive halo not host a prominent visible galaxy?
@@
This paper is organized as follows. Section 2 presents our methodology for modeling SIDM halos ($\S$ 2.1), simulating their evolutionary processes ($\S$ 2.2), and computing their lensing properties ($\S$ 2.3). In Section 3, we compare the results of our SIDM modeling with observational results ($\S$ 3.1) and evaluate the performance of gravitational lens fitting using the mock lensing image ($\S$ 3.2). Section 4 presents a discussion of our results, focusing on their implications and limitations, and finally, Section 5 summarizes the principal conclusions of this study and outlines potential directions for future research.
@@
Throughout this paper, we adopt the Planck $\rm\Lambda CDM$} cosmology (Planck2015) with $\Omega_{m}=0.307$ and $H_{0}=67.7~{\rm km}~{\rm s}^{-1}~{\rm Mpc}^{-1}$.
@@ 2.1 Solving the density profile of the SIDM halo
Considering its low computational cost and the convenience of being able to model specific galaxies and dark matter halo systems, we use a semi-analytical method called isothermal Jeans method to construct the SIDM halo. This method was first proposed by Kaplinghat2014; Kaplinghat2016 and its efficacy for SIDM density profiles has been intensively tested using hydrodynamical simulations of SIDM (e.g., Robertson2021). Here, we present the basic idea and implementation process of this method. For more detailed information, readers are referred to Jiang2023, where a public code implementation is also provided.
The model is based on the idea that energy transfer through elastic scattering between dark matter particles leads to the formation of an isothermal SIDM core after several interactions over the halo’s evolutionary timescale. The density of this isothermal core $\rho_{\rm iso}$ should be describable by the Jeans equation,
@@
where $\Phi$ is the total gravitational potential, including both dark matter and baryons. Assuming an isotropic ($\beta=0$) and constant velocity dispersion ($\nu=\nu_{0}$) in the core region, we get a simple generic solution of the Jeans equation,
@@
where $\rho_{\rm 0}$ is the central density of the isothermal core and $\Delta\Phi=\Phi(r)-\Phi(0)$. To obtain the expression of $\Phi(r)$, we apply the Poisson equation, which connects the gravitational potential to the matter density,
@@
where $\rho_{\rm b}$ represents the density distribution of the galaxy. The dark matter density (Eq. 2) and the galaxy’s density are then substituted into the Poisson equation (Eq. 3). By applying appropriate variable transformations and setting the boundary conditions, the resulting second-order differential equation is solved to obtain $\rho_{\rm iso}$ as a function of radius. More details for each step can be found in Appendix A.
For a given set of $\rho_{\rm 0}$ and $\nu_{0}$, we can determine the density profile of the isothermal core of the SIDM halo following the above steps. However, elastic scattering between dark matter particles primarily occurs in the central regions, where the particle density is sufficiently high. In the outer regions of the dark halo, the interaction between particles becomes negligible, and the dark matter distribution still follows the predictions of CDM, i.e., the NFW profile:
@@
After defining $r_{200}$ as the radius within which the mean enclosed density is 200 times the critical density $\rho_{\rm crit}$, and $M_{200}$ as the mass within $r_{200}$, the scaling parameters $\rho_{s}$ and $r_{s}$ in Eq. 4 can be replaced with the halo mass $M_{200}$ and the concentration parameter $c_{200}=r_{200}/r_{s}$. Now the SIDM halo is divided into two regions, and we need to identify the boundary that separates them. The stitching radius $r_{1}$ is defined as the radius at which the local scattering rate $\Gamma(r)$, multiplied by the halo’s lifetime $t_{\rm age}$, equals 1, marking the transition point where the behavior of the dark halo changes:
@@
Figure 1: The temporal evolution of the central density $\rho_{0}$ (left panel) and velocity dispersion $\nu_{0}$ (right panel) of an SIDM core with $M_{200}=1\times 10^{11}~{\rm M_{\odot}}$. Considering the degeneracy between the formation time $t_{\rm age}$ and the scattering cross section $\sigma_{m}$, we choose their product as the horizontal axis, with time measured in units of the cosmic time of J0946+1006 (11.06 Gyr). The blue, orange, and magenta scatter points represent the evolution for the low-density solution, high-density solution, and mirrored high-density solution, respectively. The moments corresponding to $t_{\rm merge}$ and $t_{\rm coll}$ are marked by a green dashed line and a purple dashed-dotted line, respectively. The black dots represent the moments we chose to examine the evolution of the SIDM profile, while the red star represents the moment closest to the result of Minor2021 in the 2D observational space.
Here, $\sigma_{m}$ denotes the effective self-interaction cross section per particle mass (in units of $\rm cm^{2}/g$}), which we treat as a constant following Yang_Yu2022’s framework. This approximation replaces the velocity-dependent cross section with a single characteristic value for each halo. $\rho_{\rm dm}$ is the dark matter density profile, which can be set either by the inner isothermal core or by the CDM outskirts. $\langle v_{\rm pair}(r)\rangle$ denotes the mean pairwise velocity of particles at radius $r$, and for a Maxwell–Boltzmann velocity distribution with a 1D velocity dispersion profile $\nu(r)$, it is given by $\langle v_{\rm pair}(r)\rangle=\frac{4}{\sqrt{\pi}}\nu(r)$.
@@
Once $r_{1}$ is determined, we can stitch these two regions together by imposing the matching conditions, $\rho_{\rm iso}(r_{1})\approx\rho_{\rm cdm}(r_{1})$ and $M_{\rm iso}(r_{1})\approx M_{\rm cdm}(r_{1})$, and then minimizing the “stitching error” term $\delta^{2}$:
@@
Here, we adopt the outside-in approach (Robertson2021; Sagunski2021). For a CDM halo with known $M_{200}$, $c_{200}$, and the inner galaxy’s density distribution, we first compute the NFW profile for its outskirts. For a given $t_{\rm age}\cdot\sigma_{m}$ value, we determine $r_{1}$ using Eq. 5. With a set of ($\rho_{0}$, $\nu_{0}$) values, we then derive the isothermal core density profile by solving the coupled Jeans and Poisson equations (Eq. 1 and Eq. 3). Finally, we optimize the ($\rho_{0}$, $\nu_{0}$) combination to minimize $\delta^{2}$ (Eq. 6), ensuring a smooth transition between the isothermal core and CDM outskirts at $r_{1}$. In summary, once $M_{200}$, $c_{200}$, $t_{\rm age}$, $\sigma_{m}$, and the galaxy’s density profile are provided, this isothermal Jeans model yields the complete density profile of the SIDM halo.
In Appendix A, we discuss the impact of baryonic matter on SIDM halos within our SIDM model. Given the nondetection of an optical counterpart for J0946+1006 and the upper limit of its luminosity should be no bigger than $5\times 10^{6}~{\rm L_{\odot}}$ (Vegetti_2010), we restrict our analysis exclusively to the dark-matter-only situation. Specifically, we exclude the galactic density profile from Eq. 3 and employ the NFW profile alone to derive the density profile of the inner SIDM core.